# Using a Deep Learning Model to Simulate Human Stock Traders’ Methods of Chart Analysis

### 30 Dec 2023 v2

@@ Abstract
Despite the efficient market hypothesis, many studies suggest the existence of inefficiencies in the stock market leading to the development of techniques to gain above-market returns. Systematic trading has undergone significant advances in recent decades with deep learning schemes emerging as a powerful tool for analyzing and predicting market behavior. In this paper, a method is proposed that is inspired by how professional technical analysts trade. This scheme looks at stock prices of the previous 600 days and predicts whether the stock price will rise or fall 10% or 20% within the next D days. Plus, the proposed method uses the Resnet’s (a deep learning model) skip connections and logits to increase the probability of the prediction. The model was trained and tested using historical data from both the Korean and US stock markets. We show that using the period label of 5 gives the best result. On Korea market it achieved a profit more than 49% above the market return, and a profit more than 40% above the market return on the US market.
@@ 1 Introduction
Professional technical analysts somehow make a consistent profit without these sophisticated approaches. By just looking at the charts, these human traders find major and minor trends as well as support and resistance levels of the stock. Support and resistance levels are price levels in trading where buying (support) or selling (resistance) pressure is expected to be significant, often leading to price reversals or consolidation. Inspired by these traders, this paper attempts to simulate their approach using a deep learning model. The underlying premise is that there are certain price patterns that occur before a stock rises in the short term. The model utilizes historical stock prices over a period of 600 days to predict whether there will be a 10% rise or fall in the stock price within the next D days. The Resnet model that was applied to a time series problem Wang et al. (2017) is used in this paper. Using the Resnet’s skip connections, the model is expected to capture these major and minor trends. For example, using a kernel size of 5 would extract features similar to the 5-day moving average in the first layer, and the next layer will again calculate the 5-day average with this first layer’s output which results in extracting longer-term features. When trading, the model will buy stocks that are expected to rise by 10% in the next D days. In order to increase the probability of success in stock trading, stocks whose largest softmax logit from the Resnet that exceeds a certain threshold are considered for trading. By setting this threshold, the trading is focused on stocks that the model predicts with higher confidence, potentially increasing the chances of achieving profitable trades. As far as the authors know, this paper is the first to utilize softmax logits of a deep learning model for stock trading. The model was trained using all tickers from KOSPI, KOSDAQ market for Korea stocks, and tickers from NYSE, NASDAQ, AMEX for US stocks from year 2006 to 2015, with validation period set to year 2016 to 2019. The model was tested on historical data from year 2020 to 2022, giving a profit of more than 49% above the Korea market return, and a profit of more than 40% above the US market return. Also there are lot of myth of retail technical traders, whether they can consistently make money just by looking at the charts, in contrast to large firms equipped with both financial resources and extensive information, or merely it is just a result of luck. This paper is the first to scientifically validate and cast doubt on the methods of such traders by simulating these traders approach.
@@ 2 Related Work
Using a sliding window and assigning labels to price movement with discretized return rate is not new. By discretizing the return rate, one can overcome small price change sensitivity and facilitate model training. Han et al. (2023) discussed the effect of N-Period Min Max labeling instead of just predicting whether the price will go up or down and used the XGBoost model to trade accordingly. They assigned labels to stock price movement according to which side of the upper or lower price limit of the given period is hit first. Gudelek et al. (2017) used a 2-dimensional CNN to predict the next day’s return of commonly used ETFs. 28 features with a 28-day window were used as input to the CNN which includes close price, volume, and several technical indicators such as Relative Strength Index (RSI), Simple Moving Average (SMA), Moving Average Convergence Divergence (MACD). Takeuchi & Lee (2013) used an autoencoder composed of stacked restricted Boltzmann machines. They used 12 monthly returns and 20 daily returns as features and predict whether the stock’s return will be above or below the median return of whole stocks. Mukherjee et al. (2023) trained a CNN network by inputting 2-D histograms generated out of the quantized dataset within a particular time frame. It was tested during the recent COVID-19 pandemic and achieved an accuracy of 97.66%. Our approach differs from these approaches in that only OHLCV (open, high, low, close, volume) price data was used without any human-made features, uses a long window of 600 days to capture price trends, and assign labels by whether the price hits a fixed return rate of 10% for a given period. Sample weighted loss function was also used to improve class imbalance and utilize softmax logits for better accuracy and cumulative return. Backtest results on 2020-2022 is also provided that significantly outperform the market. Providing backtest results is important because high prediction accuracy does not lead to long-term profits if the Profit and Loss (P/L) ratio is poor.
@@ 3.1 Dataset Preprocessing
OHLCV (open, high, low, close, volume) information of daily stock prices of all the tickers (KOSPI, KOSDAQ tickers in Korea market and NYSE, NASDAQ, AMEX tickers in US) are collected from Naver Finance (finance.naver.com/) for Korean stocks, and historical data from Stooq (www.stooq.com) is used for US stocks. The log price (base 10) was used instead of daily returns because we suspect the stock price range provides important information and using the original price will lead to exploding gradient. The label was determined by observing the stock price for the preceding 600 days and recording whether the stock price increases or decreases by 10% within the next D days, compared to the closing price. We believe that discretizing returns is important to facilitate learning and avoid overfitting. 10% was chosen because the Volatility Interrupt (VI), a mechanism that pauses the market for a specified period of time to allow market participants to reassess the situation and prevent excessive price swings, is activated in Korea market when the price rises by 10%. More than 10 tickers rise by 10% every day, although this number differs largely according to market conditions. Label for 20% was made also. The model would only see these OHLCV prices of 600 days (input dimension of 5*600) and would have no information of which ticker it is extracted from or any other fundamental factors such as corporate earnings, dividends, and cash flows. The model would only see the ’chart patterns’ to predict the future price. The window of 600 days is chosen to encompass the effect of 480 day moving average. In human experts’ technical analysis, the moving average is a tool used to smooth out price data over a specific time period. When the stock’s price approaches this long-term moving average (such as 480-day moving average), it can act as a support level that doesn’t let the stock price to fall below. The variable D was experimented with different values of 1, 3, 5, 10, 15, 20, and 30 to evaluate their impact on the model’s performance. In the 5-day 10% scheme, a label of 2 is given if the highest price of the next 5 days is 10% or higher, 1 if both the highest and lowest prices are in the range of -10% to 10%, and 0 if the lowest price is less than -10%, during the next 5 days. Note that the 5-day 10% scheme and 10-day 10% scheme will have the same label of 2 if the price reaches 10% on the third day and falls -10% on fourth day. If both the highest and lowest prices reached the set limit of +/-10% on the same day, they were treated as label 2 although we don’t know which limit the stock price reached first. If the lowest price reached the -10% first the trade would result in loss, while if the highest price reached +10% first it would result in gain in real trade.
@@
A well-chosen back-testing period is important to avoid over-fitting. In general, a simple buy&hold strategy is advantageous during a bullish market when prices are rising. On the other hand, in a sideways market where prices are moving within a range, it is advantageous to trade when the price touches either end of the range. Strategies optimized specifically for either upside (bullish) or sideways markets may not work as effectively in different market conditions. Hence training and testing periods were carefully chosen to avoid this. The training period was chosen to be 2006-2015, and the validation period to be 2016-2019, both of which include strong bear and bull markets and markets that go sideways. The model was tested from 2020 to 2022. For US dataset, tickers that have close prices above \$2000 and below \$2 were excluded in training, as stock prices range widely from penny stocks to above \$50,$0$. Korea stock dataset consists of 2673488 training samples, 1611189 validation samples, and 1477200 test samples while US stock dataset consists of 3807869 training samples, 2409155 validation samples, and 2369288 test samples.
@@
Dataset statistics for Korea and US stocks are shown in table 1 and 2. Each item in the table shows a percentage of 10% or 20% rise, fall, and the stocks that went sideways in this range during the given period rounded to second decimal separated by ’/’. For example, the training dataset of US with labeling period 3 shows 0.07/0.07/0.86, which means that during the period of 3 days, roughly 7% of stocks reached a 10% price rise, 7% of stocks reached 10% price fall, and 86% stayed in the price range of -10% 10% fall during 2006-2015. The ’uncertain’ columns shows the percentage of the samples classified as 10% rise although we don’t know because the price fluctuated twice the given label percentage. Because the uncertain percentage is at most 13%, it would be reasonable to label them as 10% rise. However, to adopt a more conservative stance during backtesting, these instances are treated as losses. It can be seen that the longer the labeling period, the 10% rise or fall ratio gets higher and the sideways ratio gets lower. This can be interpreted as stocks that move sideways determining direction over time.
@@ 3.2 Proposed Model
ResNet He et al. (2016) is a deep learning model architecture designed to overcome the vanishing gradient problem by incorporating skip connections. These connections enable the network to retain important information from earlier layers, making it easier to train deeper neural networks. Wang et al. (2017) modified Resnet’s architecture to be used in time series. This model is modified for training. The Resnet architecture’s skip connections is expected to capture major and minor trends, lower layers to capture short-term trends, and higher layers the long-term. For example, using a kernel size of 5 would extract features similar to the 5-day moving average in the first layer, and the next layer will again calculate the 5-day average with this first layer’s output which results in extracting longer-term features. This process continues through the subsequent layers. And because skip connection is used, this information would be gradually integrated. The architecture of Wang et al. (2017) is modified in this application. The architecture of Wang et al. (2017) and the model prposed in this paper is shown in 1 and 2. The model gets an input dimension of 5x600, representing the log OHLCV price of 600 stock days. A one-dimensional Resnet model with 5 blocks with a fixed channel number of 12 is utilized, then passed to the Global Average Pooling (GAP) layer to get the mean values of each feature map. Each block consists of 3 convolutional layers with kernel sizes of 7,5,3 each with Batch Normalization (BN) and ReLu activation applied. A skip connection is added after each block. The number of channels is fixed to 12 all thorough. After passing the Fully Connected Layer softmax values were calculated representing the probability of a 10% (20%) rise, 10% (20%) fall, and price between -10% to 10% (-230% to 20%). The Binary Cross Entropy (BCE) is used as the loss function. When implementing the model, Resnet implementation for time series in this paper Wang et al. (2017) is used. THhe model is trained with kernel stride of 1, padding of 1. All models were trained for 50 epochs, with 16-bit precision. Training took 1.5 days for Korea stocks and 5 days for US stocks. The same model using varying kernel sizes 5,15,30,60,120 (denoted as TrendUp) and 5,15,30,60,120 (denoted as TrendDown) was also trained. The kernel size is same within block with thses models. These varying kernel sizes were expected to capture various short and long-term trends but were shown to be ineffective compared to fixed kernel size of 7,5,3.
3.3 Validation Accuracy and Thresholding
@@
val_us_accuracy_result.csv
Table 3: Validation set statistics for US label
[…]
val_kr_accuracy_result.csv
Table 4: Validation set statistics for Korea label
The softmax logit (output) of a model is commonly believed to indicate the level of confidence in the result of an image recognition task, such as identifying 90% happiness and 5% surprise in emotion recognition. One might wonder if this value really represents the certainty of the result. This research indicates that applying the concept of thresholding to trading domains can lead to increased confidence in predictions and higher probabilities of success. By setting a specific threshold value for the softmax output, it is possible to enhance cumulative profit. This phenomenon, which we refer to as ’thresholding’, has been validated in our study. Only samples with the largest softmax value of the last layer of the model exceeding a certain threshold are considered when calculating accuracy.
The validation results which calculates accuracy, f1 macro score, precision and recall for each label name, threshold, and regarding dataset proportion is shown in table 4 and 3. All numbers are rounded up to four decimals. Threshold of 0, 0.7, 0.8, 0.9, 0.99, 0.999, 0.9995 are tested. Each statistics are calculated only with the items that are output by model above the given threshold. ’dataset proportion’ denotes what fraction of its original dataset the metrics were calculated. Precision is the ratio tp / (tp + fp) where tp is the number of true positives and fp the number of false positives. Recall is the ratio tp / (tp + fn) which is a measure of the ability of a classification model to capture and correctly identify all relevant instances. ’f1 macro’ denotes the F1 score with macro averaging. It calculates F1 score for each label and averages them. F1 score is the harmonic mean of precision and recall. ’nan’ represents not a number since no instances were outputted to calculate those scores. It can be seen that as you increase the threshold, the overall accuracy increases while the corresponding proportion of dataset decreases. Because the model only predicts confident instances, the accuracy reaches almost 1 (100%) for some labels with dataset proportion close to 0 (0%). Some models seems to fail in the learning process, specializing in predicting specific labels. For example, US model label_1_tp20_ls20 only predict for label 1 above threshold 0.8, which gives 1 in label 1 precision but all others 0. In US model label_20_tp20_ls20, the precision of label 1 increase as you increase the threshold but precision of label 0 and 2 decreases as you increase the threshold. We rank the models according to f1 macro scores and and take the threshold which gives the highest score for each labels. Out of these we proceed to backtesting with which its dataset proportion is above 0.000001 to ensure not to few trades are made.
We rank models by their F1 macro scores and choose the threshold that maximizes the score for each label. Subsequently, we only take the label and threshold which the respective dataset proportion exceeds 0.00001 to ensure an adequate number of trades are made. The according label name with respective threshold, dataset proportion, f1 score, and the accuracy ordered by f1 score is shown on table 6 for Korea market5 for US market. We proceed them to backtesting.
@@
3.4 Validaton Backtest Results
@@
Table 7: Korea backtest result on validation period
@@
Table 8: US backtest result on validation period
@@
The models trained by chosen labels above is backtested on from year 2016 to 2019. Original starting cash was set to be 10 million won for Korea stocks and ten thousand dollars for US stocks. Trading commissions were set to be 0.015% every time trade is made and tax were set to be 0.20%, only applied when the trade ended with a profit. We assume that the whole amount can be sold when the price reaches the given percentage, hence only +/-10% of P&L (Profit and Loss) will occur during single trade. The entry ratio was set to be 0.1, which represents entering every trade 1/10 of original cash. We show the average result of five runs. The results are shown on table 7 for Korea market and table 8 order by total return. After the labeling period, a postfix ’_sidecut_True’ is assigned if the stock was sold at the closing price on days when it went sideways during the labeling period. Alternatively, postfix ’_sidecut_False’ is assigned if the stock was held even after the labeling period until it reached a given percentage of +/-10%. For example, label_5_tp10_ls10_threshold_0.99_sidecut_True refers to the model trained for a labeling period of 5 days, backtested with threshold 0.99, with entry ratio 0.1, and sells at the close price if the stock went sideways during the labeling period. In the real trading scenario, a 10% drop followed by a 10% increase will result in a loss, and vice versa as a gain. Because minute data for all tickers are not available in this research, it is unclear whether profit was made if the price fluctuated twice the labeling percentage on that day. For comparison, we present the results of holding randomly selected stocks until they hit a predetermined percentage limit over the given period. These are shown with prefix ’random_’ in front of the label. Despite the stocks are randomly selected, some number of backtests outperform the market return, which is 14.54%, -1.17%, for KOSPI, KOSDAQ (Korea market) and 60.52%, 83.0%, 20.33% for S&P500, NASDAQ, NYSE, AMEX (US market). This maybe due to survival bias, in which we backtested with the companies that survived until 2023, and high turnover rate, which has the effect similar to investing higher amount of initial capital because when a stock is sold another stock is purchased immediately with the available cash. ’Total Trades’ shows the total number of trades executed within the given period. ’Win Rate’ shows the percentage of profitable trades out of the total number of trades. Labels with ’_s’ has lower win rate compared to ’_w’ because the model has to predict stock prices either rise, falls, or remain within the given percentage range in the given period. The ’Win/Loss Ratio’ represents the number of winning trades to losing trades, excluding scenarios where the stock price goes sideways in the given percentage range. ’Maximum Drawdown’ is the maximum percentage decline in the trading account from its peak value to its lowest point. It is a measure that indicates largest unrealized loss the investor has to suffer during the trading period. Sharpe Ratio is a risk-adjusted measure of return, the average return earned in excess of the risk-free rate per unit of volatility. We notice that although label_1_tp10_ls10 gives the highest F1 macro score for both Korea and Us labels, it doesn’t given necessarily gives the best total return. This seems to be do to over-fitting, as there may be no underlying reasons that makes us possible to predict the stocks that will rise 10% in that day and the model just fits to noise signals.

### 08 Apr 2024 v3

@@ Abstract
Despite the efficient market hypothesis, many studies suggest the existence of inefficiencies in the stock market leading to the development of techniques to gain above-market returns. Systematic trading has undergone significant advances in recent decades with deep learning schemes emerging as a powerful tool for analyzing and predicting market behavior. In this paper, a method is proposed that is inspired by how professional technical analysts trade. This scheme looks at stock prices of the previous 600 days and predicts whether the stock price will rise or fall 10% or 20% within the next D days. The proposed method uses the Resnet’s (a deep learning model) skip connections and logits to increase the probability of the prediction. The model was trained and tested using historical data from both the Korea and US stock markets. The backtest is done using the data from 2020 to 2022. Using the proposed method for the Korea market it gave return of 75.36% having Sharpe ratio of 1.57, which far exceeds the market return by 36% and 0.61, respectively. On the US market it gives total return of 27.17% with Sharpe ratio of 0.61, which outperforms other benchmarks such as NASDAQ, S&P500, DOW JONES index by 17.69% and 0.27, respectively.
@@ 1 Introduction
Professional technical analysts somehow make a consistent profit without these sophisticated approaches. By just looking at the charts, these human traders find major and minor trends as well as support and resistance levels of the stock. Support and resistance levels are price levels in trading where buying (support) or selling (resistance) pressure is expected to be significant, often leading to price reversals or consolidation. Inspired by these traders, this paper attempts to simulate their approach using a deep learning model. The underlying premise is that there are certain price patterns that occur before a stock rises in the short term. The model utilizes historical stock prices over a period of 600 days to predict whether there will be a 10% rise or fall in the stock price within the next D days. We assume that there are specific spanning time that this patterns occur, and try to find them. The Resnet model that was applied to a time series problem Wang et al. (2017) is used in this paper. Although one might suggest LSTM is more appropriate for time series problems, we think Resnet is suit for the task because these traders trade by drawing trend lines on charts (which is a 2D image). For these traders, finding out short-term and long-term trends is important when trading. Using the Resnet’s skip connections, the model is expected to capture these major and minor trends. For example, using a kernel size of 5 would extract features similar to the 5-day moving average in the first layer, and the next layer will again calculate the 5-day average with this first layer’s output which results in extracting longer-term features. When trading, the model will buy stocks that are expected to rise by 10% or 20% in the next D days. We view this as classification problem instead of time series regression because in stock market there are instances where a stock may rise but the timing of its rise is uncertain. In order to increase the probability of success in stock trading, stocks whose largest softmax logit from the Resnet that exceeds a certain threshold are considered for trading. By setting this threshold, the trading is focused on stocks that the model predicts with higher confidence, potentially increasing the chances of achieving profitable trades. As far as the authors know, this paper is the first to utilize softmax logits of a deep learning model for stock trading. The model was trained using all tickers from KOSPI, KOSDAQ market for Korea stocks, and tickers from NYSE, NASDAQ, AMEX for US stocks from year 2006 to 2015, with validation period set to year 2016 to 2019. The model was tested on historical data from year 2020 to 2022, giving a profit of more than 49% above the Korea market return, and a profit of more than 40% above the US market return.
@@ 2 Related Work
Using a sliding window and assigning labels to price movement with discretized return rate is not new. By discretizing the return rate, one can overcome small price change sensitivity and facilitate model training. Han et al. (2023) discussed the effect of N-Period Min Max labeling instead of just predicting whether the price will go up or down and used the XGBoost model to trade accordingly. They assigned labels to stock price movement according to which side of the upper or lower price limit of the given period is hit first. Teixeira & de Oliveira (2010) combined technical indicators with nearest neighbor classification on fifteen real stocks from Sao Paulo Stock Exchange. Since trading volume provides important information, Chavarnakul & Enke (2008) combines equivolume charting with generalized regression neural network to predict S&P 500 index. Dash & Dash (2016) proposed computational efficient functional link artificial neural network (CEFLANN) with technical indicators to output trading signal for BSE SENSEX and S&P 500 index. Gudelek et al. (2017) used a 2-dimensional CNN to predict the next day’s return of commonly used ETFs. 28 features with a 28-day window were used as input to the CNN which includes close price, volume, and several technical indicators such as Relative Strength Index (RSI), Simple Moving Average (SMA), Moving Average Convergence Divergence (MACD). Takeuchi & Lee (2013) used an autoencoder composed of stacked restricted Boltzmann machines. They used 12 monthly returns and 20 daily returns as features and predict whether the stock’s return will be above or below the median return of whole stocks. Mukherjee et al. (2023) trained a CNN network by inputting 2-D histograms generated out of the quantized dataset within a particular time frame. It was tested during the recent COVID-19 pandemic and achieved an accuracy of 97.66%. Our approach differs from these approaches in that only OHLCV (open, high, low, close, volume) price data was used without any human-made features, uses a long window of 600 days to capture price trends, and assign labels by whether the price hits a fixed return rate of 10% for a given period. Sample weighted loss function was also used to improve class imbalance and utilize softmax logits for better accuracy and cumulative return. Backtest results on 2020-2022 is also provided that significantly outperform the market. Providing backtest results is important because high prediction accuracy does not lead to long-term profits if the Profit and Loss (P/L) ratio is poor.
@@ 3.1 Dataset Preprocessing
OHLCV (open, high, low, close, volume) information of daily stock prices of all the tickers (KOSPI, KOSDAQ tickers in Korea market and NYSE, NASDAQ, AMEX tickers in US) are collected from Naver Finance (finance.naver.com/) for Korea stocks, and historical data from Stooq (www.stooq.com) is used for US stocks. The log price (base 10) was used instead of daily returns because we suspect the stock price range provides important information and using the original price will lead to exploding gradient. The label was determined by observing the stock price for the preceding 600 days and recording whether the stock price increases or decreases by 10% within the next D days, compared to the closing price. We believe that discretizing returns is important to facilitate learning and avoid overfitting. 10% was chosen because the Volatility Interrupt (VI), a mechanism that pauses the market for a specified period of time to allow market participants to reassess the situation and prevent excessive price swings, is activated in Korea market when the price rises by 10%. More than 10 tickers rise by 10% every day, although this number differs largely according to market conditions. Label for 20% was made also. The model would only see these OHLCV prices of 600 days (input dimension of 5*600) and would have no information of which ticker it is extracted from or any other fundamental factors such as corporate earnings, dividends, and cash flows. The model would only see the ’chart patterns’ to predict the future price. The window of 600 days is chosen to encompass the effect of 480 day moving average. In human experts’ technical analysis, the moving average is a tool used to smooth out price data over a specific time period. When the stock’s price approaches this long-term moving average (such as 480-day moving average), it can act as a support level that doesn’t let the stock price to fall below. The variable D was experimented with different values of 1, 3, 5, 10, 15, 20, and 30 to evaluate their impact on the model’s performance. In the 5-day 10% scheme, a label of 2 is given if the highest price of the next 5 days is 10% or higher, 1 if both the highest and lowest prices are in the range of -10% to 10%, and 0 if the lowest price is less than -10%, during the next 5 days. Note that the 5-day 10% scheme and 10-day 10% scheme will have the same label of 2 if the price reaches 10% on the third day and falls -10% on fourth day. If both the highest and lowest prices reached the set limit of +/-10% on the same day, they were treated as label 2 although we don’t know which limit the stock price reached first. If the lowest price reached the -10% first the trade would result in loss, while if the highest price reached +10% first it would result in gain in real trade.
@@
A well-chosen back-testing period is important to avoid over-fitting. In general, a simple buy&hold strategy is advantageous during a bullish market (when stocks prices tend to keep increasing). On the other hand, in a sideways market where prices are moving within a range, it is advantageous to trade when the price touches either end of the range. Strategies optimized specifically for either upside or sideways markets may not work as effectively in different market conditions. Hence training and testing periods were carefully chosen to avoid this. The training period was chosen to be 2006-2015, and the validation period to be 2016-2019, both of which include strong bear and bull markets and markets that go sideways. The model was tested from 2020 to 2022. For US dataset, tickers that have close prices above \$2000 and below \$2 were excluded in training, as stock prices range widely from penny stocks to above \$50,$0$. Korea stock dataset consists of 2673488 training samples, 1611189 validation samples, and 1477200 test samples while US stock dataset consists of 3807869 training samples, 2409155 validation samples, and 2369288 test samples.
@@
Dataset statistics for Korea and US stocks are shown in table 1 and 2. Each item in the table shows a percentage of 10% or 20% rise, fall, and the stocks that went sideways in this range during the given period rounded to second decimal separated by ’/’. For example, the training dataset of US with labeling period 3 shows 0.07/0.07/0.86, which means that during the period of 3 days, roughly 7% of stocks reached a 10% price rise, 7% of stocks reached 10% price fall, and 86% stayed in the price range of -10% 10% fall during 2006-2015. The ’uncertain’ columns shows the percentage of the samples classified as 10% rise although we don’t know because the price fluctuated twice the given label percentage. However, to adopt a more conservative stance during backtesting, these instances are treated as losses. It can be seen that the longer the labeling period, the 10% rise or fall ratio gets higher and the sideways ratio gets lower. This can be interpreted as stocks that move sideways determining direction over time.
@@ 3.2 Proposed Model
ResNet He et al. (2016) is a deep learning model architecture designed to overcome the vanishing gradient problem by incorporating skip connections. These connections enable the network to retain important information from earlier layers, making it easier to train deeper neural networks. Wang et al. (2017) modified Resnet’s architecture to be used in time series. This model is modified for training. The Resnet architecture’s skip connections is expected to capture major and minor trends, lower layers to capture short-term trends, and higher layers the long-term. For example, using a kernel size of 5 would extract features similar to the 5-day moving average in the first layer, and the next layer will again calculate the 5-day average with this first layer’s output which results in extracting longer-term features. This process continues through the subsequent layers. And because skip connection is used, this information would be gradually integrated. The architecture of Wang et al. (2017) is modified in this application. The architecture of Wang et al. (2017) and the model proposed in this paper is shown in 1 and 2. The model gets an input dimension of 5x600, representing the log OHLCV price of 600 stock days. A one-dimensional Resnet model with 5 blocks with a fixed channel number of 12 is utilized, then passed to the Global Average Pooling (GAP) layer to get the mean values of each feature map. Each block consists of 3 convolutional layers with kernel sizes of 7,5,3 each with Batch Normalization (BN) and ReLu activation applied. A skip connection is added after each block. The number of channels is fixed to 12 all thorough. After passing the Fully Connected Layer softmax values were calculated representing the probability of a 10% (20%) rise, 10% (20%) fall, and price between -10% to 10% (-230% to 20%). The Binary Cross Entropy (BCE) is used as the loss function. When implementing the model, Resnet implementation for time series in this paper Wang et al. (2017) is used. The model is trained with kernel stride of 1, padding of 1. All models were trained for 50 epochs, with 16-bit precision. Training took 1.5 days for Korea stocks and 5 days for US stocks. The same model using varying kernel sizes 5,15,30,60,120 and 5,15,30,60,120 was also trained. The kernel size is same within block with these models. These varying kernel sizes were expected to capture various short and long-term trends but were shown to be ineffective compared to fixed kernel size of 7,5,3.
3.3 Choosing the Best Model and Threshold
The softmax logit (output) of a model is commonly believed to indicate the level of confidence in the result of an image recognition task, such as identifying 90% happiness and 5% surprise in emotion recognition. We believe this is also the case for trading. By only predicingt for which the largest softmax value of the model’s last layer (logit) is above a certain threshold, we can increase model profitability. We refer this as ‘threshodling’. This would not be possible for normal classification problem as we have to give result for all the instances, but in trading we can only confident stocks.
We show part of the validation result table for model trained on US market with various thresholds on table 3. Accuracy, f1 macro score, precision and recall for each label name, threshold, and regarding dataset proportion is shown. All numbers are rounded up to four decimals. Threshold of 0, 0.7, 0.8, 0.9, 0.99, 0.999, 0.9995 are tested. Each statistics are calculated only with the items that are output by model above the given threshold. ’dataset proportion’ denotes what fraction of its original dataset the metrics were calculated. Precision is the ratio tp / (tp + fp) where tp is the number of true positives and fp the number of false positives. Recall is the ratio tp / (tp + fn) which is a measure of the ability of a classification model to capture and correctly identify all relevant instances. ’f1 macro’ denotes the F1 score with macro averaging. It calculates F1 score for each label and averages them. F1 score is the harmonic mean of precision and recall. ’nan’ represents not a number since no instances were outputted to calculate those scores.
@@ 3.3 Choosing the Best Model and Threshold
val_us_accuracy_result_selected.csv
Table 3: Part of Validation set statistics for US label
It can be seen that as you increase the threshold, the overall accuracy increases while the corresponding proportion of dataset decreases. Because the model only predicts confident instances, the accuracy reaches almost 1 (100%) for some labels with dataset proportion close to 0 (0%). Some models seems to fail in the learning process, specializing in predicting specific labels. For example, US model label_1_tp20_ls20 only predict for label 1 above threshold 0.8, which gives 1 in label 1 precision but all others 0. In US model label_20_tp20_ls20, the precision of label 1 increase as you increase the threshold but precision of label 0 and 2 decreases as you increase the threshold. We think this problem arises due to dataset imbalance.
Selecting the appropriate label and threshold for backtesting is a problem since testing multiple combinations will only expose us to multiple comparison bias. To choose the best model, we hypothesize that successful model should be able to sufficiently predict all three labels (up, down, sideways) to learn the important features. We rank models by their F1 macro scores and choose the threshold that maximizes the score for each label. Subsequently, we only take the label and threshold which the respective dataset proportion exceeds 0.00001 to ensure an adequate number of trades are made. The according label name with respective threshold, dataset proportion, f1 score, and the accuracy ordered by f1 score is shown on table 5 for Korea market4 for US market. We proceed to backtesting with these labels.
@@
When using the model for trading, the model looks all the tickers at the end of the market every day, and buys the stock if the model predicts label 1 and the respective softmax logit is above threshold. Stocks bought are automatically sold if it increases or decreases 10% or 20%, or sold if N days passed according to trained label. The stock buying process using the model is illustrated in algorithm 11
Algorithm 1 Stock selecting scheme using trained model
1:  Initialize model parameters and threshold
2:  for each trading day do
3:     for each ticker do
4:        Retrieve 600 OHLCV price data for the ticker
5:        Feed data into the model
6:        Obtain predicted class and largest logit
7:        if largest logit $>$ threshold AND predicted class = 1 then
8:           Execute buy order for the stock
9:        end if
10:     end for
11:  end for
In order to choose which model with the given period label performs the best, backtesting is conducted during the validation period. Original starting cash was set to be 10 million won for Korea stocks and ten thousand dollars for US stocks. Trading commissions were set to be 0.015% every time trade is made and tax were set to be 0.20% which is applied whenever the trade ended with a profit. We assume that the whole amount can be sold when the price reaches the given percentage, hence only +/-10% or +/-20% of P&L (Profit and Loss) will occur during single trade. The entry ratio was chosen to be 1/20 of available cash. Although other entry scheme such as using fixed entry amount of 1/10 of original cash (100\$ per trade for US) and double it when the original cash doubles, this scheme was chosen to reduce the variance of each runs and compare with the index. Top ten results when ordered by Sharpe ratio is shown on table 6 and 7.
@@
Table 6: Korea market validation period backtest result top 10 ordered by Sharpe ratio
@@
Table 7: US market validation period backtest result top 10 ordered by Sharpe ratio
@@
After the labeling period, a postfix ’_sidecut_True’ is assigned if the stock was sold at the closing price on days when it went sideways during the labeling period. Alternatively, postfix ’_sidecut_False’ is assigned if the stock was held even after the labeling period until it reached a given percentage of +/-10%. For example, label_5_tp10_ls10_threshold_0.99_sidecut_True refers to the model trained for a labeling period of 5 days, backtested with threshold 0.99, with entry ratio 0.1, and sells at the close price if the stock went sideways during the labeling period. In the real trading scenario, a 10% drop followed by a 10% increase will result in a loss, and vice versa as a gain. Because minute data for all tickers are not available in this research, it is unclear whether profit was made if the price fluctuated twice the labeling percentage on that day. ’Trades’ shows the total number of trades executed within the given period. Buying and selling one ticker counts as one trade. ’Return’ shows the total return generated with respect to original cash during the test period. ’MDD’ represents maximum draw-down is the maximum percentage decline in the trading account from its peak value to its lowest point. It is a measure that indicates largest unrealized loss the investor has to suffer during the trading period. ’Sharpe Ratio’ is a risk-adjusted measure of return, the average return earned in excess of the risk-free rate per unit of volatility. The shapre ratio is used to compare various strategies since total return can be magnified with increased volatility using leverage. Higher volatility also means higher loss as well as higher profit. We choose our best model to be model_label_20_tp10_ls10_sidecut_False for Korea market and model_label_30_tp20_ls20_sidecut_False for US market, which ranks the best far exceeding others by 0.36 and 0.16 on Sharpe ratio.
To ascertain the impact of threshold adjustments on model returns, we conducted backtesting on our top-performing model across a range of thresholds. As the threshold increases, both accuracy and F1 macro score tend to increase, but the proportion of the dataset that the model predicts decreases. We examined thresholds ranging from 0 (no threshold applied) to 0.9995. For each threshold, we executed five runs and analyzed the resulting Sharpe ratios. The boxplots illustrating these Sharpe ratios during the validation period for KR market is shown on Figure 3 and and US market on Figure 4. Threshold affects how frequently model will buy stocks as well as model accuracy. Because of this, we expected the Sharpe ratio plotted against the threshold to resemble a mountain shape, with the peak representing the optimal balance between accuracy and coverage. However for Korea market this was the case, for US market the Sharpe ratio tend to decrease as you increase the threshold. We believe this difference is likely due to market conditions. During the 4-year validation period, the Korean stock market only saw a maximum increase of 14.54%, whereas the US S&P 500 and NASDAQ rose by 60.52% and 83% respectively. In a strongly bullish market like the US during this period, holding a wider range of stocks including even less certain ones, would be advantageous. Even uncertain stocks are likely to end in profit.
Figure 3: Box Plot of Sharpe Ratio with different thresholds during validation period using KR Best Model
Figure 4: Box Plot of Sharpe Ratio with different thresholds during validation period using US Best Model